Question-5: Solve the system of equations by Gaussian elimination: 24 +32 = 8 2.1 + 3y + 2 = 5 x - y - 22 = – 5
Answer #1

Au niven that 24+ 32 = 8 2x+3y +2 = 5 a-9-22=-5 Argumented mebix is so 2318 1) Interchage Ist and 3rd equation) R, Rz 2 3 1 xNow, dividing equation Tird by 2. R3 I R3. a y -22=-5 r -1-2 yt z = 3 4 + 32=4 Now, Add-1 times the equation Ind to the equat